OpenStudy (anonymous):

Hii =) help?? solve the following 2x-y=11 4x+3y=17

OpenStudy (anonymous):

x = 5, y = -1

OpenStudy (anonymous):

multiply first equation by 3 and then add. the y's will drop out

OpenStudy (anonymous):

the first thing you need to do is solve for one of the two variables. since the top one already has y by itself lets solve that one for y. we get: y=2x-11 now that we have an equivalent for y, we can plug that into the second equation. it will look like this: 4x + 3(2x-11)=17 now we can solve for x: 4x+6x-33=17 10x-33=17 10x=50 x=5 now that we know x, we can plug that back into our equation for y to get the y value: y=2(5)-11 y=10-11 y=-1 so the answer is x=5, y=-1

OpenStudy (anonymous):

Thankyou all very thoroughly explained naper =)

OpenStudy (anonymous):

lol, it's just how i do. ;) you are very welcome.

OpenStudy (anonymous):

2x-y=11 4x+3(2x-11) y=2(3.3)+11 -y=-2x+11/-1 4x+6x-33 y=6.6+11 y=2x-11 10x=33/10 y=17.6 x=3.3